OpenStudy (anonymous):

A spacecraft that is initially at rest turns on its engine at t = 0. If its mass is m = 2600 kg and the force from the engine is 40 N, what is the acceleration of the spacecraft?

OpenStudy (anonymous):

Newton's 2nd Law: F = m * a

OpenStudy (shamim):

40=2600*a a=?